### v2

@@ Abstract
Process mining offers powerful techniques for discovering, analyzing, and enhancing real-world business processes. In this context, Petri nets provide an expressive means of modeling process behavior. However, directly analyzing and comparing intricate Petri net presents challenges. This study introduces PetriNet2Vec, a novel unsupervised methodology based on Natural Language Processing concepts inspired by Doc2Vec and designed to facilitate the effective comparison, clustering, and classification of process models represented as embedding vectors. These embedding vectors allow us to quantify similarities and relationships between different process models. Our methodology was experimentally validated using the PDC Dataset, featuring 96 diverse Petri net models. We performed cluster analysis, created UMAP visualizations, and trained a decision tree to provide compelling evidence for the capability of PetriNet2Vec to discern meaningful patterns and relationships among process models and their constituent tasks. Through a series of experiments, we demonstrated that PetriNet2Vec was capable of learning the structure of Petri nets, as well as the main properties used to simulate the process models of our dataset. Furthermore, our results showcase the utility of the learned embeddings in two crucial downstream tasks within process mining enhancement: process classification and process retrieval.
@@
Process Mining Model enhancement Embedding vectors Petri Nets
@@ 1 Introduction
Modern business process models exhibit a level of complexity that renders traditional analysis tools insufficient for comprehensive understanding and optimization [20]. This challenge demands innovative approaches capable of extracting complex relationships governing process behavior.
Process discovery techniques leverage event data to extract, analyze, and visualize the actual execution of business processes. Generally, a data-driven approach with advanced algorithms is used to construct models, such as Petri nets, that accurately capture sequences of activities, i.e. transitions (activity), and places (state) [16]. Petri nets are mathematical and visual modeling tools used to describe distributed systems. These networks allow us to assess process performance, identify bottlenecks, and uncover hidden patterns through conformance-checking techniques. Ultimately, facilitating process enhancement and optimization.
Traditional process mining core capabilities often fall short when faced with the scale and complexity of modern processes [7]. They may struggle to handle the vast amounts of data generated and lack the sophistication to uncover complex patterns that machine learning techniques could reveal [3].
Embedding vectors are numerical representations of objects or concepts in a continuous vector space [9]. Commonly used in Natural Language Processing (NLP) related tasks, which often involve machine learning algorithms. Embedding vectors capture semantic relationships between entities. In process mining, embeddings can provide a powerful means of representing complex process structures and relationships within process models [19]. For example, we can embed individual activities, control flow structures, or entire process models in vector representations. This enables sophisticated similarity analysis, identification of analogous patterns, and application of predictive modeling techniques that would be difficult to achieve with traditional process representations.
Our proposed approach enables us to encode both the structural information of Process models, in Petri net format, and the individual tasks into compact vector representations (i.e. embedding vectors), facilitating various downstream tasks such as similarity analysis and process discovery.
@@ 2 Related works
Process comparison is crucial in conformance analysis, process enhancement, knowledge transfer, and process retrieval. Existing techniques for process comparison broadly fall into three categories: behavioral analysis, structural analysis, and task comparison [22]. Behavioral strategies focus on the order of activities within execution logs, while structural approaches analyze the process model as a graph. Task comparison offers a granular view, looking at relationships between individual activities [4].
Existing research on process model similarity analysis offers valuable tools, but certain limitations demand further attention. Behavioral methods can lack effectiveness when variations in terminology mask similarities between processes [6]. This issue is particularly pronounced in localized comparison methods, where direct node-to-node comparisons fail to grasp similarities between models from different domains. Additionally, both behavioral and structural techniques frequently encounter exponential complexity stemming from concurrency and loops within process models [6]. Furthermore, structural approaches reliant on graph edit distance face computational scalability challenges with large graphs due to their NP-Hard complexity [28].
@@
Behavioral process comparison focuses on the execution sequences observed in event logs. Van Der Aalst et al. [24] emphasize the value of direct log analysis in uncovering process behavior. Dijkman et al. [4] introduce the concept of “causal footprints”, which capture the temporal dependencies between activities to represent process behavior. Kunze et al. [10] highlight relationships between exclusivity and strict order as key behavioral aspects. Van Dongen et al. [25] develop behavioral metrics for quantifying process similarities and differences. Sánchez-Charles et al. [23] extend this by proposing a behavioral distance measure to compare process behavior holistically.
@@
Alternatively, structural methods examine the overall process model as a graph, analyzing elements like node relationships, paths, connectivity, and control flow. Standard techniques utilize edit graph distance to calculate the similarity between process graphs [5]. More nuanced methods consider richer information, such as node types (e.g., AND/XOR) and sequential information [15]. Zhou et al. [29] enhance structural techniques by incorporating insights from execution logs and weighting graph edges based on their frequency. While valuable for analyzing overall process flow, such techniques may not fully capture fine-grained behavioral variations that arise from the diverse execution of process activities.
@@
Task comparison provides a detailed analysis of activity relationships within process models. Node-by-node techniques excel at identifying subtle differences, while block-based methods reveal more significant structural changes [12, 4, 1, 27]. Recent clustering approaches hold the potential for comparing localized process sections [18]. However, these localized clustering methods may struggle when analyzing processes with highly variable flows or behavioral patterns.
@@
In the domain of Similarity-based retrieval of semantic graphs, Hoffman et al. [8] proposed an approach employing Graph Neural Networks (GNNs) for process mining, framing it as a graph matching challenge. Despite the inherent complexity of GNNs, necessitating larger datasets for effective training, their methodology shows significant promise, particularly in retrieval scenarios essential to Process-Oriented Case-Based Reasoning (POCBR) within smart manufacturing - a focus distinct from our current study. Nonetheless, in Section5.4, we comprehensively evaluate our methodology, specifically in the realm of semantic retrieval tasks for process models, utilizing PetriNet2Vec. This assessment underscores the resilience and adaptability of our retrieval approach, highlighting its potential for broader applications in the field.
@@
The optimal comparison strategy hinges on the desired analytical depth and specific research questions. While current techniques offer valuable insights, their limitations suggest the need for a method that offers a more holistic view of process execution, considering both behavioral patterns and the broader process structure. Our proposed approach addresses these needs. Firstly, it considers the entire process and its structure, mirroring structural approaches. Secondly, it enables the comparison of multiple processes and clusters, echoing the strengths of most behavioral approaches.
@@ 3.1 Learning embeddings with doc2vec and graph2vec
The doc2vec methodology [11], an extension of the word2vec concept, has become popular for enabling the simultaneous learning of embedding vectors for both documents and individual words. In word2vec [14], a word is predicted based on the words in its context. Figure 1 illustrates the functioning of the CBOW (Continuous Bag-of-Words) approach. While this figure is useful for illustrating the word2vec concept, we can reinterpret the optimization objective function during training as a binary classifier, predicting $P(1|w_{i},w_{c})$ when the word $w_{i}$ is part of the word context $w_{c}=\{\dots,w_{i-2},w_{i-1},w_{i+1},w_{i+2},\dots\}$. Thus, if $w_{i}$ is not in the context $w_{c}$, then we have $P(0|w_{i},w_{c})$. Then, during training, the negative sampling strategy generates positive and negative example pairs $(w_{i},w_{c})$.
@@
The graph2vec algorithm is a powerful formulation that allows us to learn to embed vectors for graph representation [17]. Given a graph $G_{n}=(V,N)$ consisting of a set of vertices $V$ and edges $N$, graph2vec naturally extends the principles of doc2vec. By analogy, each node in the graph can be seen as a word $V_{i}\equiv w_{i}$, with the neighboring nodes $V_{c}\equiv w_{c}$ directly connected to $V_{i}$, serving as its context. Consequently, if each node is associated with a token $t_{k}$, as depicted in Figure 1c, we can concurrently learn embeddings for the entire graph $h_{n}$ and for each individual node.
This approach outperforms other graph embedding techniques because it trains on various graph samples and incorporates neighborhood information for every node. Thus, it helps the neural network learn complex graph structures more effectively. As a result, graph2vec embeddings can accurately capture similarities between graphs with similar structures, making them highly useful across different downstream applications [2].
@@ 3.2 Cluster algorithm
As discussed in the previous sections, one of the objectives of this work is to learn embedding for each Petri net model and group them by similarity. For this purpose, we adopted a version of the Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) clustering algorithm with cosine distance [13].
@@ 3.3 Process models dataset
Our research leverages the PDC Dataset [26], specifically curated for the Process Discovery Contest. This dataset encompasses a total of 96 Petri net models stored in PNML format. As a specialized class of models, Petri nets present a versatile range of configurations, rendering them indispensable for evaluating process discovery algorithms and techniques.
@@ 4 Methodology for Learning Petri Net Embeddings
To train PetriNet2Vec, we represent a model as the set of all possible transition pairs within it. Furthermore, we adopted the Distributed Memory (DM) algorithm implemented by the Gensim [21] library in Python, as illustrated on the right side of Figure 3. The diversity of process models in the training dataset ensures that some pairs of transitions are more frequent in some process models than others, and some transition pairs may even exist in some models but not in others. Therefore, by training using the described tuples, we can ensure that the learned vectors represent every model uniquely, carrying information about their structure. During training, we aim to maximize the probability of task $t_{i+1}$ being in the context of task $t_{i}$ and model $m_{j}$. We train on the entire dataset, maximizing the average probability loss:
@@
This training method does not require any model labels or supervision; in other words, this embedding training is unsupervised. To improve the separability between the learned vectors, we adopted the negative sampling strategy [14]. With this approach, in every training epoch, some pairs are randomly sampled from other models to form negative tuples for training that do not exist in the current model being trained. With negative sampling, the maximization objective becomes $P(1|t_{i+1},t_{i},m_{j})$, i.e., predict 1 when the task $t_{i+1}$ is in the context of task $t_{i}$ and model $m_{j}$. Nevertheless, all these sampling and training epochs occur in parallel for speed and efficiency improvements.
(a) Methodoly 1 (‘None’)
(b) Methodoly 2
@@ 5.1 Cluster analysis
As previously discussed, specific tasks within our Petri net models may be unspecified, denoted by black boxes. We face two choices in developing our methodology: we can adopt a generic and uniform name for all black boxes in all models, for instance, using the ‘None’ token, akin to an out-of-vocabulary token in word2vec, or uniquely label each black box according to its position in the network,, referred to as ‘nx’ transitions from Figure 4. These alternatives for constructing the task dictionary (also known as the dictionary of tokens) influence the embedding vectors we aim to learn during PetriNet2Vec model training. Thus, our initial inquiry delves into how the quality of the clusters is impacted under each scenario.
@@
Table 1 compares the silhouette scores for the two methodologies. The leftmost column denotes the size of embeddings $d_{n}$, while the adjacent columns vary the hyperparameter named ‘minimum cluster size’ within the HDBSCAN clustering algorithm. Lower minimum cluster sizes elevate the probability of forming smaller clusters, whereas higher values tend to yield larger, amalgamated clusters. Given the silhouette score’s sensitivity to cluster size, caution is warranted in adjusting this parameter beyond these bounds. After numerous iterations, we identified an acceptable range between 2 and 4.
Silhouette scores were computed by running each methodology ten times, with 1000 training epochs for each investigated embedding size. The central values in the table represent the averages across these runs, while values in parentheses are standard deviations. Consistently, Methodology 1 yields superior silhouette scores compared to Methodology 2, indicating more cohesive clusters. Optimal results were attained with $d_{n}=8$ and a minimum cluster size of four. However, slight overlaps in standard deviations across the $d_{n}=8$ row suggest that any minimum cluster size, i.e., 2, 3, or 4, could be considered viable for grouping PNML models using HDBSCAN and cosine distance. Henceforth, we will adhere to and explore Methodology 1, employing embeddings of size $d_{n}=8$ and a minimum cluster size of 4.
Table 1: Comparison of silhouette scores between Methodology 1 and Methodology 2. Silhouette scores were averaged across ten runs with 1000 training epochs for each embedding size. Methodology 1 consistently yields superior results, particularly evident with an embedding size of $d_{n}=8$.
Methodology 1 (‘None’) Methodology 2
$d_{n}$ 2 3 4 2 3 4
4 0.61 (0.10) 0.69 (0.04) 0.69 (0.04) 0.56 (0.05) 0.62 (0.04) 0.51 (0.07)
8 0.71 (0.02) 0.72 (0.02) 0.73 (0.01) 0.48 (0.03) 0.45 (0.02) 0.25 (0.03)
16 0.47 (0.01) 0.45 (0.03) 0.38 (0.03) 0.32 (0.02) 0.32 (0.02) 0.21 (0.02)
32 0.42 (0.02) 0.41 (0.00) 0.34 (0.02) 0.29 (0.01) 0.28 (0.01) 0.17 (0.01)

### v3

@@ Abstract
Process Mining offers a powerful framework for uncovering, analyzing, and optimizing real-world business processes. Petri nets provide a versatile means of modeling process behavior. However, traditional methods often struggle to effectively compare complex Petri nets, hindering their potential for process enhancement. To address this challenge, we introduce PetriNet2Vec, an unsupervised methodology inspired by Doc2Vec. This approach converts Petri nets into embedding vectors, facilitating the comparison, clustering, and classification of process models. We validated our approach using the PDC Dataset, comprising 96 diverse Petri net models. The results demonstrate that PetriNet2Vec effectively captures the structural properties of process models, enabling accurate process classification and efficient process retrieval. Specifically, our findings highlight the utility of the learned embeddings in two key downstream tasks: process classification and process retrieval. In process classification, the embeddings allowed for accurate categorization of process models based on their structural properties. In process retrieval, the embeddings enabled efficient retrieval of similar process models using cosine distance. These results demonstrate the potential of PetriNet2Vec to significantly enhance process mining capabilities.
@@
Process Mining , Model Clustering , Embedding Vectors , Petri Nets
journal: Intelligent Systems with Applications
Highlights
  1. 1.
    A Novel Approach for Process Model Embeddings: This research introduces PetriNet2Vec, a new method inspired by Natural Language Processing (NLP) to create embedding vectors for process models represented as Petri nets.
  2. 2.
    Learning from Sentences, Understanding Processes: PetriNet2Vec treats process paths like sentences in documents, allowing it to leverage NLP techniques for process mining tasks.
  3. 3.
    Unsupervised Learning for Robust Embeddings: PetriNet2Vec learns effectively through unsupervised training with negative sampling, capturing the intricacies of various process models.
  4. 4.
    Empowering Downstream Tasks: The learned embeddings prove valuable for process classification and retrieval, which are essential tasks in process mining.
  5. 5.
    Real-World Applications: This study demonstrates success with a simulated dataset and proposes future work involving real-world data and incorporating broader task context for improved process discovery.
@@ 1 Introduction
Modern business process models exhibit a level of complexity that renders traditional analysis tools insufficient for comprehensive understanding and optimization [Recker et al., 2009]. This challenge demands innovative approaches capable of extracting complex relationships governing process behavior.
Process discovery techniques leverage event data to extract, analyze, and visualize the actual execution of business processes. Generally, a data-driven approach with advanced algorithms is used to construct models, such as Petri nets, that accurately capture sequences of activities, i.e. transitions (activity), and places (state) [Murata, 1989]. Petri nets are mathematical and visual modeling tools used to describe distributed systems. These networks allow us to assess process performance, identify bottlenecks, and uncover hidden patterns through conformance-checking techniques. Ultimately, facilitating process enhancement and optimization.
Traditional process mining core capabilities often fall short when faced with the scale and complexity of modern processes [Ganesh et al., 2023]. They may struggle to handle the vast amounts of data generated and lack the sophistication to uncover complex patterns that machine learning techniques could reveal [Cárdenas Maita et al., 2018].
Embedding vectors are numerical representations of objects or concepts in a continuous vector space [Jurafsky & Martin, ]. Commonly used in Natural Language Processing (NLP) related tasks, which often involve machine learning algorithms. Embedding vectors capture semantic relationships between entities. In process mining, embeddings can provide a powerful means of representing complex process structures and relationships within process models [Pismerov & Pikalov, 2022]. For example, we can embed individual activities, control flow structures, or entire process models in vector representations. This enables sophisticated similarity analysis, identification of analogous patterns, and application of predictive modeling techniques that would be difficult to achieve with traditional process representations.
@@ 1.1 Problem statement
1.2 Contributions and overview
This work introduces PetriNet2Vec, a novel method inspired by NLP to create embedding vectors for process models represented as Petri nets. By treating process paths like sentences in documents, PetriNet2Vec leverages NLP techniques to capture the intricacies of various process models through unsupervised training with negative sampling. Our approach encodes both the structural information of process models in Petri net format and the individual tasks into compact vector representations (i.e., embedding vectors), facilitating downstream tasks such as similarity analysis, process retrieval, and classification of process properties. Demonstrated on a simulated dataset, this approach shows promise for real-world applications, with future work aimed at incorporating broader task contexts to enhance process discovery. Additionally, we have made the PetriNet2Vec Python package publicly available for the community, which can be installed via the Python Package Index (PyPI)111pip install PetriNet2Vec.
The article is organized as follows: Section 2 provides a comprehensive review of related works in process mining, highlighting the existing methods, unveiling their limitations, and comparing them with our proposed PetriNet2Vec method through applying embedding techniques. In Section 3, we present the theoretical background of learning vector embeddings (e.g. word2vec), followed by the clustering algorithm and dataset description. Section 4 details our methodology for learning Petri net embeddings, including the transformation of Petri nets into intermediate representations and the training process of PetriNet2Vec. Section 5 presents the results of our experiments, including cluster analysis, visual assessments, and examples of downstream tasks such as process classification and retrieval. Finally, Section 6 concludes the manuscript with a summary of our findings, limitations, and suggestions for future research directions.
@@ 2 Related works
Process comparison is crucial in conformance analysis, process enhancement, knowledge transfer, and process retrieval. Existing techniques for process comparison broadly fall into three categories: behavioral analysis, structural analysis, and task comparison [Syukriilah et al., 2015]. Behavioral strategies focus on the order of activities within execution logs, while structural approaches analyze the process model as a graph. Task comparison offers a granular view, looking at relationships between individual activities [Dijkman et al., 2011].
2.1 Behavioral analysis
@@
Behavioral process comparison focuses on the execution sequences observed in event logs. Van Der Aalst et al. [2006] emphasize the value of direct log analysis in uncovering process behavior. These methods, while effective in extracting sequences of activities, face challenges with terminology variations. PetriNet2Vec embeds entire process models, effectively capturing structural dependencies and semantic similarities, thus reducing the impact of terminology variations and improving scalability in handling complex processes.  Dijkman et al. [2011] introduce the concept of “causal footprints”, which capture the temporal dependencies between activities to represent process behavior. Kunze et al. [2011] highlight relationships between exclusivity and strict order as key behavioral aspects. Van Dongen et al. [2013] develop behavioral metrics for quantifying process similarities and differences. Sánchez-Charles et al. [2016] extend this by proposing a behavioral distance measure to compare process behavior holistically.
Although useful, those methods can be limited by the complexity introduced by concurrency and loops [Dijkman et al., 2013]. PetriNet2Vec’s embedding approach captures both dependencies and concurrency more efficiently, providing a comprehensive representation that scales better with complex processes.
2.2 Structural analysis
@@
Alternatively, structural methods examine the overall process model as a graph, analyzing elements like node relationships, paths, connectivity, and control flow. Standard techniques utilize edit graph distance to calculate the similarity between process graphs [Dijkman et al., 2009]. More nuanced methods consider richer information, such as node types (e.g., AND/XOR) and sequential information [Montani et al., 2015]. While these methods are effective approaches reliant on graph edit distance, they face computational scalability challenges with large graphs due to their NP-Hard complexity [Zeng et al., 2009]. By representing process models as embedding vectors, PetriNet2Vec overcomes scalability issues and encodes structural information more efficiently, allowing for faster and more accurate comparisons.
Zhou et al. [2019] enhance structural techniques by incorporating insights from execution logs and weighting graph edges based on their frequency. While valuable for analyzing overall process flow, such techniques may not fully capture fine-grained behavioral variations that arise from the diverse execution of process activities. PetriNet2Vec integrates both structural and behavioral details in its embeddings, enabling finer-grained analysis and capturing nuances that purely structural methods might miss.
2.3 Task comparison
@@
Task comparison provides a detailed analysis of activity relationships within process models. Node-by-node techniques excel at identifying subtle differences, while block-based methods reveal more significant structural changes [Li et al., 2008, Dijkman et al., 2011, Bae et al., 2006, Yan et al., 2012]. On the other side, PetriNet2Vec provides a unified embedding representation, facilitating both detailed node-level and holistic structural comparisons, offering a more balanced approach to process model comparison.
Recent clustering approaches hold the potential for comparing localized process sections [Peeva & van der Aalst, 2023]. However, these localized clustering methods may struggle when analyzing processes with highly variable flows or behavioral patterns but can struggle with highly complex and diverse process models. By learning embeddings for tasks and entire processes, PetriNet2Vec enhances clustering capabilities, effectively managing variability and complexity in process models.
@@
In the domain of Similarity-based retrieval of semantic graphs, Hoffmann & Bergmann [2022] proposed an approach employing Graph Neural Networks (GNNs) for process mining, framing it as a graph matching challenge. Despite the inherent complexity of GNNs, necessitating larger datasets for effective training, their methodology shows significant promise, particularly in retrieval scenarios essential to Process-Oriented Case-Based Reasoning (POCBR) within smart manufacturing - a focus distinct from our current study. Nonetheless, in Section5.4, we comprehensively evaluate our methodology, specifically in the realm of semantic retrieval tasks for process models, utilizing PetriNet2Vec. This assessment underscores the resilience and adaptability of our retrieval approach, highlighting its potential for broader applications in the field.
2.4 Conclusions on Related Work
@@
While current techniques offer valuable insights, their limitations highlight the need for a method that provides a more holistic view of process execution, considering both behavioral patterns and the broader process structure. Our proposed approach effectively addresses these needs. It considers the entire process and its structure, mirroring structural approaches, and enables the comparison of multiple processes and clusters, capturing the strengths of behavioral approaches.
PetriNet2Vec’s comprehensive embedding approach offers a unified representation, facilitating both detailed node-level and task sequence modeling. This dual capability enhances the analysis and retrieval of complex process models, overcoming limitations of existing methods and demonstrating higher accuracy and scalability in clustering and classification tasks.
@@ 3.1 Learning embeddings with doc2vec and graph2vec
The doc2vec methodology [Le & Mikolov, 2014], an extension of the word2vec concept, has become popular for enabling the simultaneous learning of embedding vectors for both documents and individual words. In word2vec [Mikolov et al., 2013], a word is predicted based on the words in its context. Figure 1 illustrates the functioning of the CBOW (Continuous Bag-of-Words) approach. While this figure is useful for illustrating the word2vec concept, we can reinterpret the optimization objective function during training as a binary classifier, predicting $P(1|w_{i},w_{c})$ when the word $w_{i}$ is part of the word context $w_{c}=\{\dots,w_{i-2},w_{i-1},w_{i+1},w_{i+2},\dots\}$. Thus, if $w_{i}$ is not in the context $w_{c}$, then we have $P(0|w_{i},w_{c})$. Then, during training, the negative sampling strategy generates positive and negative example pairs $(w_{i},w_{c})$.
@@
The graph2vec algorithm is a powerful formulation that allows us to learn to embed vectors for graph representation [Narayanan et al., 2017]. Given a graph $G_{n}=(V,N)$ consisting of a set of vertices $V$ and edges $N$, graph2vec naturally extends the principles of doc2vec. By analogy, each node in the graph can be seen as a word $V_{i}\equiv w_{i}$, with the neighboring nodes $V_{c}\equiv w_{c}$ directly connected to $V_{i}$, serving as its context. Consequently, if each node is associated with a token $t_{k}$, as depicted in Figure 1c, we can concurrently learn embeddings for the entire graph $h_{n}$ and for each individual node.
This approach outperforms other graph embedding techniques because it trains on various graph samples and incorporates neighborhood information for every node. Thus, it helps the neural network learn complex graph structures more effectively. As a result, graph2vec embeddings can accurately capture similarities between graphs with similar structures, making them highly useful across different downstream applications [Cai et al., 2018].
@@ 3.2 Cluster algorithm
As discussed in the previous sections, one of the objectives of this work is to learn embedding for each Petri net model and group them by similarity. For this purpose, we adopted a version of the Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) clustering algorithm with cosine distance [McInnes et al., 2017].
@@ 3.3 Process models dataset
Our research leverages the PDC Dataset [Verbeek, 2023], specifically curated for the Process Discovery Contest. This dataset encompasses a total of 96 Petri net models stored in PNML format. As a specialized class of models, Petri nets present a versatile range of configurations, rendering them indispensable for evaluating process discovery algorithms and techniques.
@@ 4 Methodology for Learning Petri Net Embeddings
To train PetriNet2Vec, we represent a model as the set of all possible transition pairs within it. Furthermore, we adopted the Distributed Memory (DM) algorithm implemented by the Gensim [Řehůřek & Sojka, 2010] library in Python, as illustrated on the right side of Figure 3. The diversity of process models in the training dataset ensures that some pairs of transitions are more frequent in some process models than others, and some transition pairs may even exist in some models but not in others. Therefore, by training using the described tuples, we can ensure that the learned vectors represent every model uniquely, carrying information about their structure. During training, we aim to maximize the probability of task $t_{i+1}$ being in the context of task $t_{i}$ and model $m_{j}$. We train on the entire dataset, maximizing the average probability loss:
@@
This training method does not require any model labels or supervision; in other words, this embedding training is unsupervised. To improve the separability between the learned vectors, we adopted the negative sampling strategy [Mikolov et al., 2013]. With this approach, in every training epoch, some pairs are randomly sampled from other models to form negative tuples for training that do not exist in the current model being trained. With negative sampling, the maximization objective becomes $P(1|t_{i+1},t_{i},m_{j})$, i.e., predict 1 when the task $t_{i+1}$ is in the context of task $t_{i}$ and model $m_{j}$. Nevertheless, all these sampling and training epochs occur in parallel for speed and efficiency improvements.
(a) Methodoly 1 (‘None’).
(b) Methodoly 2.
@@ 5.1 Cluster analysis
As previously discussed, specific tasks within our Petri net models may be unspecified, denoted by black boxes. We face two choices in developing our methodology: we can adopt a generic and uniform name for all black boxes in all models, for instance, using the ‘None’ token, akin to an out-of-vocabulary token in word2vec, or uniquely label each black box according to its position in the network, referred to as ‘nx’ transitions from Figure 4. These alternatives for constructing the task dictionary (also known as the dictionary of tokens) influence the embedding vectors we aim to learn during PetriNet2Vec model training. Thus, our initial inquiry delves into how the quality of the clusters is impacted under each scenario.
@@
Tables 1 and 2 compare the silhouette scores for the different methodologies. The leftmost column indicates the size of embeddings, $d_{n}$, while the adjacent columns vary the only hyperparameter, ‘minimum cluster size’, within the HDBSCAN clustering algorithm using cosine distance. Lower minimum cluster sizes increase the likelihood of forming smaller clusters, whereas higher values tend to produce larger, amalgamated clusters. Since the silhouette score is sensitive to cluster size, it is crucial to adjust this parameter within reasonable bounds. We determined that an acceptable range is between 2 and 6, as these values balance the need for distinct, well-defined clusters with the necessity of maintaining meaningful groupings, making them effective for process retrieval and clustering.
The silhouette scores were computed by running each methodology 30 times, with 500 training epochs for each investigated embedding size. The central values in the table represent the averages across these runs, while the values in parentheses indicate the standard deviations. Consistently, Methodology 1 (Table 1) yielded superior silhouette scores compared to Methodology 2 (Table 2), indicating more cohesive clusters. Optimal results were attained with $d_{n}=[4,8]$ and a minimum cluster size of $[3,4,5,6]$. Although $d_{8}$ does not have the highest score, it demonstrates the best compromise by achieving minimal variance, indicating more consistent cluster formation across the 30 runs.
We applied the t-test with a significance level of $p=0.01$ for two independent samples to determine if there is a tie between the mean silhouette scores of the groups. This test evaluates whether the observed difference between the means is statistically significant, allowing us to detect ties between hyperparameter combinations. The null hypothesis states that the means of the two groups are equal, while the alternative hypothesis posits that the means are different. We tested the scores resulting from $d_{8}$ and a minimum cluster size of 4 against all other pairs of hyperparameter combinations. The bold values in Table 1 indicate where the null hypothesis was accepted, showing no significant difference in means. Consequently, we decided to explore Methodology 1, employing embeddings of size $d_{8}$ and a minimum cluster size of 4.
Table 1: Comparison of silhouette scores between Methodology 1 (‘None’). Silhouette scores were averaged across 30 runs with 500 training epochs for each embedding size. Methodology 1 consistently yields superior results, particularly evident with embedding sizes of $d_{n}=[4,8]$.
Minimum cluster size
$d_{n}$ 2 3 4 5 6
4 0.64 (0.09) 0.71 (0.07) 0.74 (0.07) 0.74 (0.07) 0.74 (0.06)
8 0.7 (0.02) 0.71 (0.02) 0.72 (0.02) 0.73 (0.04) 0.73 (0.03)
16 0.48 (0.01) 0.48 (0.02) 0.37 (0.04) 0.36 (0.02) 0.37 (0.01)
32 0.46 (0.01) 0.4 (0.03) 0.33 (0.02) 0.35 (0.01) 0.35 (0.01)
Table 2: Comparison of Silhouette Scores between Methodology 2. The silhouette scores were averaged across 30 runs with 500 training epochs for each embedding size. Methodology 2 showed worse results compared to Methodology 1.
Minimum cluster size
$d_{n}$ 2 3 4 5 6
4 0.58 (0.07) 0.62 (0.08) 0.48 (0.08) 0.51 (0.07) 0.54 (0.05)
8 0.44 (0.04) 0.45 (0.04) 0.27 (0.05) 0.26 (0.06) 0.32 (0.04)
16 0.32 (0.02) 0.31 (0.02) 0.2 (0.02) 0.19 (0.02) 0.17 (0.07)
32 0.3 (0.01) 0.29 (0.01) 0.18 (0.01) 0.17 (0.03) 0.17 (0.01)